When 12 moles of carbon monoxide reacts with 9 moles of oxygen to produce carbon dioxide in a closed 10 liter container. what will be the total pressure in the flask at 0 degree Celsius after the reaction?
1
Expert's answer
2016-03-09T08:36:42-0500
2CO + O2 = 2CO2
O2 in excess, all calculations will be provided on CO2 amount
reacted N(CO) = 12; N(O2) = 6 moles
formed N(CO2) = 12 moles
resedue N(O2)=9-6=3 moles
In flask after reaction 12 + 3 = 15 moles of gases.
PV=NRT P=NRT/V
P=15*8,314*273/10= 3 404,6 KPa